If P20,000 will accumulate to P45,758.55 in 4 years, what is the interest rate compounded semi-annually?

1
Expert's answer
2021-12-29T16:04:59-0500

"A=P(1+\frac {r} {100}) ^n"


"45,758.55=20,000(1+\frac {r} {100}) ^8"


"\frac {45,758.55} {20,000} =(1+\frac {r} {100}) ^8"


"2.2879 =(1+\frac {r} {100}) ^8"


"\sqrt[8] {2.2879} =\sqrt[8] {(1+\frac {r} {100}) ^8}"


"1.109=1+\frac {r} {100}"


"0.109=\frac {r} {100}"


"r=10.9\%×2"


r=21.8%